# Goal-Conditioned Data Augmentation for Offline Reinforcement Learning

### v1

@@ Abstract
Offline reinforcement learning (RL) enables policy learning from pre-collected offline datasets, relaxing the need to interact directly with the environment. However, limited by the quality of offline datasets, it generally fails to learn well-qualified policies in suboptimal datasets. To address datasets with insufficient optimal demonstrations, we introduce Goal-cOnditioned Data Augmentation (GODA), a novel goal-conditioned diffusion-based method for augmenting samples with higher quality. Leveraging recent advancements in generative modeling, GODA incorporates a novel return-oriented goal condition with various selection mechanisms. Specifically, we introduce a controllable scaling technique to provide enhanced return-based guidance during data sampling. GODA learns a comprehensive distribution representation of the original offline datasets while generating new data with selectively higher-return goals, thereby maximizing the utility of limited optimal demonstrations. Furthermore, we propose a novel adaptive gated conditioning method for processing noised inputs and conditions, enhancing the capture of goal-oriented guidance. We conduct experiments on the D4RL benchmark and real-world challenges, specifically traffic signal control (TSC) tasks, to demonstrate GODA’s effectiveness in enhancing data quality and superior performance compared to state-of-the-art data augmentation methods across various offline RL algorithms.
Index Terms:
Offline reinforcement learning, diffusion model, data augmentation
I Introduction
@@
Reinforcement learning [1] aims to learn a control policy from trial and error through interacting with the environment. While demonstrating remarkable performance in various domains, this approach typically requires vast amounts of training data collected from these interactions. Such data-intensive requirements become impractical in applications where environmental interactions are costly, risky, or time-consuming, such as robotics [2, 3], autonomous driving [4, 5] and TSC [6, 7]. Offline RL offers a feasible solution to these challenges by enabling policy learning directly from pre-collected historical datasets, thus significantly reducing the need to interact directly with the environment.
@@
Although offline RL makes policy learning less expensive, its performance is highly dependent on the quality of the pre-collected datasets and may suffer from lack of diversity, behavior policy bias, distributional shift, and suboptimal demonstrations [8]. The performance of offline RL tends to decline drastically when trained with suboptimal offline datasets. Previous studies have attempted to address these issues by constraining the learned policy to align closely with the behavior policy [9] or by limiting out-of-distribution action values [10]. Although these approaches have shown performance improvements, they retain the inherent defects of offline datasets, remaining highly dependent on data quality.
@@
Several studies have addressed the limitations of offline RL using data augmentation methods to generate more diverse samples. One approach involves learning world models to mimic environmental dynamics and iteratively generate synthetic rollouts from a start state [11, 12]. While this method significantly improves sample efficiency and data diversity, it suffers from compounding errors and fails to control the quality of generated trajectories. Other research leverages generative models to capture the distributions of collected datasets and randomly sample new transition data [13]. Although these methods demonstrate some performance improvements, they remain inefficient when dealing with datasets containing limited optimal demonstrations. This inefficiency stems from their inability to effectively control the quality of generated data.
@@
We attempt to address this challenge by taking advantage of generative modeling to augment higher-quality data with directional goals. Unlike previous studies that sample data unconditionally and randomly [13], we introduce GODA to incorporate representative goals, guiding the samples toward higher returns. Given the exceptional performance of diffusion models [14, 15] in the field of generative artificial intelligence, GODA utilizes a diffusion model as its generative framework. GODA is trained to capture a comprehensive representation of the data distribution from the original dataset while sampling new data conditioned on selective high-return goals. This approach maximizes the utility of the limited well-performed trajectories in the original datasets.
@@
Inspired by Decision Transformer [16], we define the ’goal’ as the return-to-go (RTG), which represents the cumulative rewards from the current step until the end, coupled with its specific timestep in trajectories. RTG explicitly indicates the expected future rewards for a given behavior at the current timestep for a specific trajectory. We assume that at the same timestep across different trajectories, a higher RTG signifies a higher goal. To generate samples that exceed the quality of the original dataset, we introduce three goal selection mechanisms and a scaling technique to control our expected goals during sampling.
@@
3) We show GODA’s competence through comprehensive experiments on the D4RL benchmark compared with state-of-the-art data augmentation methods across multiple offline RL algorithms. We further evaluate the effectiveness of GODA on a real-world application, i.e., traffic signal control with small-size datasets obtained from widely used controllers in real-world deployments. These evaluations verify GODA’s effectiveness in addressing various challenges, significantly enhancing the applicability of RL-based methods for real-world scenarios.
@@
The remaining sections are organized as follows: in Section II, we review recent researches on offline RL and data augmentation. Some preliminaries about offline RL and diffusion models are introduced in Section III, followed by details about our methodology in Section IV. Section V describes the experimental settings about D4RL and TSC tasks, baselines and evaluation algorithms. Next, we show the data quality measurement for our augmented datasets in Section VI and provide more detailed experimental results in Section VII. Finally, we conclude our research and discuss future directions in Section VIII.
II Related Work
II-A Offline Reinforcement Learning
Offline RL involves learning policies from pre-collected offline datasets comprising trajectory rollouts generated by behavior policies. This approach is promising because it avoids the costs and risks associated with direct interactions with the environment. Conventional offline RL methods aim to alleviate the distributional shift problem, i.e., a significant drop in performance due to deviations between learned policy and the behavior policy used for generating the offline data [17]. To address this issue, various strategies have been employed, including explicit correction [18], such as constraining the policy to a restricted action space [19], and making conservative estimates of the value function [20, 21], with the aim of aligning the behavior policy with the learned policy.
Some recent studies exploit the strong sequence modeling ability of Transformer models to solve offline RL with trajectory optimization. For instance, Decision Transformer [16] and its variants [22, 23] utilize a GPT model to autoregressive predict actions given the recent subtrajectories composed of historical RTGs, states, and actions. These approaches integrate hindsight return information, i.e., RTG, with sequence modeling, eliminating the necessity for dynamic programming.
Diffusion models have also been adopted in offline RL given its exceptional capability of multi-modal distribution modeling. Diffuser [24] employs diffusion models for long-horizon planning, effectively bypassing the compounding errors associated with classical model-based RL. Hierarchical Diffuser [25] enhances this approach by introducing a hierarchical structure, specifically a jumpy planning method, to improve planning effectiveness further.
II-B Data Augmentation in Offline Reinforcement Learning
Rather than passively reusing data and concentrating on enhancing learning algorithms, data augmentation proactively generates more diverse data to improve policy optimization. Some model-based RL methods employ learned world models from offline datasets to simulate the environment and iteratively generate synthetic trajectories, facilitating policy optimization [11]. For instance, TATU [11] uses world models to produce synthetic trajectories and truncates those with high accumulated uncertainty. However, model-based RL often suffers from compounding errors in the learned world models. Some basic data augmentation functions (DAFs), i.e., translate, rotate, and reflect, are also applied to augment trajectory segments [26]. GuDA [27] further introduces human guidance into these DAFs to improve data quality, while human intervention is costly and lacks scalability. Diffusion models are also directly applied to data augmentation through the sampling process. SynthER [13] is the first work that employs diffusion models to learn the distribution of initial offline datasets and unconditionally augment large amounts of new random data. However, it fails to control the sampling process to steer toward high-return directions actively. DiffStitch [4] attempts to enhance the quality of generated data by actively connecting low-reward trajectories to high-reward ones using a stitching technique.
Figure 2: The denoiser neural network with adaptive gated conditioning architecture. The condition-adaptive gated long skip connection, shown in the left part, captures both low- and high-level features, assigning varying importance weights to each. The adaptive gated residual, depicted in the right part, further enhances the model by selectively preserving input information based on the given conditions.
III Preliminaries
III-A Offline Reinforcement Learning
@@
In RL, the task environment is generally formulated as a Markov decision process (MDP) $\{\mathcal{S},\mathcal{A},\mathcal{R},\mathcal{P},\gamma\}$ [1]. $s\in\mathcal{S}$, $s^{\prime}\in\mathcal{S}$, $a\in\mathcal{A}$, $r=\mathcal{R}(s,a)$, $\mathcal{P}(s^{\prime}|s,a)$, and $\gamma\in[0,1)$ represent state, next state, action, reward function, state transition, and discount factor, respectively. RL aims to train an agent to interact with the environment and learn a policy $\pi$ from experience. The objective of RL is to maximize the expected discounted cumulative rewards over time:
@@
III-B Diffusion Models
Diffusion models [28, 14, 15], a class of well-known generative modeling methods, aim to learn a comprehensive representation of the data distribution $p_{\text{data}}(\mathbf{x}^{N})$ with a standard deviation $\sigma_{\text{data}}$ from a given dataset. Diffusion models generally have two primary processes, the forward process, also known as the diffusion process, and the reverse/sampling process.
@@ III-B Diffusion Models
For reverse process, Denoising Diffusion Probabilistic Model (DDPM) [14] models it as a Markov chain that involves denoising an initial noise $p\left(\mathbf{x}^{0}\right)=\mathcal{N}\left(\mathbf{x}^{0};\mathbf{0},\mathbf{I}\right)$ to the original data distribution with learned Gaussian transitions. Elucidated Diffusion Model (EDM) [15] formulates the forward and reverse processes as a probability-flow ODE, where the noise level can be increased or decreased by moving the ODE forward or backward in time:
@@
where $\dot{\sigma}(t_{i})$ denotes derivative over denoise time and $\nabla_{\mathbf{x}}\log p(\mathbf{x};\sigma(t_{i}))$ is referred to as the score function [29], which points towards regions of higher data density. It is worth noting that we use $t_{i}$ to denote the noise time to distinguish it from the trajectory timestep $t$. The ODE pushes the samples away from the data or closer to the data through infinitesimal forward or backward steps. The corresponding step sequence is $\left\{t_{0},t_{1},...t_{N}\right\}$, where $t_{N}=0$ and $N$ denotes the number of ODE solver iterations.
EDM proposes to estimate the score function using denoising score matching [15]. Specifically, a denoiser neural network $D_{\theta}(\mathbf{x};\sigma)$ is trained to approximate data $\mathbf{x}^{N}$ sampled from $p_{\text{data}}$ by minimizing the $L_{2}$ denoising loss independently for each $\sigma$:
@@
Subsequently, the score function can be calculated as $\nabla_{\mathbf{x}}\log p(\mathbf{x};\sigma)=(D_{\theta}(\mathbf{x};\sigma)-\mathbf{x})/\sigma^{2}$.
IV Methodology
In this section, we introduce GODA, a goal-conditioned data augmentation method utilizing generative modeling for augmenting higher-quality synthetic transition data. Our adopted diffusion model learns comprehensive data distribution from the initial offline dataset, subsequently sampling new data towards higher return with controllable selective goal conditions. In this part, we define a representative goal for GODA and introduce different selective mechanisms for goal conditions. To proactively control the sampling direction, a controllable goal scaling factor is introduced. For better integrating goal conditions as guidance within the diffusion model, we further propose a novel adaptive gated conditioning approach that introduces condition-adaptive gate mechanism into long skip connection and residual connection.
IV-A Return-oriented goal
Prior diffusion-based work [13] lacks the ability to guide the sampling process in the desired direction. We attempt to introduce a return-oriented goal as a condition of our diffusion model. Inspired by Decision Transformer, we adopt RTG [16], cumulative rewards from the current step till the end, as the explicit goal condition
@@ IV-A Return-oriented goal
For each transition sample represented as a tuple $(s,a,s^{\prime},r)$ within a trajectory, RTG quantifies the expected future rewards for the current behavior, effectively serving as a goal. In other words, a higher RTG at a specific timestep typically signifies a higher goal for the policy to pursue. Since the same behavior at different timesteps often yields varying RTGs across different trajectories, we combine the RTG with its corresponding timestep in the trajectory as the condition for each specific transition sample. The timestep signal acts as a timestamp for each goal.
IV-B Selective goal conditions
During dataset preprocessing, we first organize offline samples into trajectories, compute the RTG for each, and append timesteps to every sample. To fully leverage well-performing samples and augment samples with higher returns, we propose three distinct condition selection mechanisms: return-prior, RTG-prior, and random goal conditions.
(1) Return-prior goal condition. In this approach, we rank all trajectories based on their return values and select the top $n$ trajectories. During the sampling process of the diffusion model, the RTG and timestep pairs $(\hat{g}_{t},t)$ from these top $n$ trajectories are selected as the sampling goal conditions. This method filters high-return trajectories from the initial offline datasets and reuses them to sample more well-optimized transitions.
(2) RTG-prior goal condition. We group RTGs by their associated timesteps and then sort them to select the top $n$ RTGs along with their corresponding timesteps as goal conditions. This approach selectively reuses high-RTG transitions for data augmentation, focusing on transitions that are most likely to yield higher returns.

### v2

@@ Abstract
Offline reinforcement learning (RL) enables policy learning from pre-collected offline datasets, relaxing the need to interact directly with the environment. However, limited by the quality of offline datasets, it generally fails to learn well-qualified policies in suboptimal datasets. To address datasets with insufficient optimal demonstrations, we introduce Goal-cOnditioned Data Augmentation (GODA), a novel goal-conditioned diffusion-based method for augmenting samples with higher quality. Leveraging recent advancements in generative modelling, GODA incorporates a novel return-oriented goal condition with various selection mechanisms. Specifically, we introduce a controllable scaling technique to provide enhanced return-based guidance during data sampling. GODA learns a comprehensive distribution representation of the original offline datasets while generating new data with selectively higher-return goals, thereby maximizing the utility of limited optimal demonstrations. Furthermore, we propose a novel adaptive gated conditioning method for processing noisy inputs and conditions, enhancing the capture of goal-oriented guidance. We conduct experiments on the D4RL benchmark and real-world challenges, specifically traffic signal control (TSC) tasks, to demonstrate GODA’s effectiveness in enhancing data quality and superior performance compared to state-of-the-art data augmentation methods across various offline RL algorithms.
1 Introduction
@@
Reinforcement learning (Sutton & Barto, 2018) aims to learn a control policy from trial and error through interacting with the environment. While demonstrating remarkable performance in various domains, this approach typically requires vast amounts of training data collected from these interactions. Such data-intensive requirements become impractical in applications where environmental interactions are costly, risky, or time-consuming, such as robotics (Bhateja et al., 2024; Tang et al., 2025), autonomous driving (Li et al., 2024; Taghavifar et al., 2025), and TSC (Zhai et al., 2025; Du et al., 2024). Offline RL offers a feasible solution to these challenges by enabling policy learning directly from pre-collected historical datasets, thus significantly reducing the need to interact directly with the environment.
@@
Although offline RL makes policy learning less expensive, its performance is highly dependent on the quality of the pre-collected datasets and may suffer from a lack of diversity, behavior policy bias, distributional shift, and suboptimal demonstrations (Prudencio et al., 2023). The performance of offline RL tends to decline drastically when trained with suboptimal offline datasets. Previous studies have attempted to address these issues by constraining the learned policy to align closely with the behavior policy (Lyu et al., 2022) or by limiting out-of-distribution action values (Kostrikov et al., 2021). Although these approaches have shown performance improvements, they retain the inherent defects of offline datasets, remaining highly dependent on data quality.
@@
Several studies have addressed the limitations of offline RL using data augmentation methods to generate more diverse samples. One approach involves learning world models to mimic environmental dynamics and iteratively generate synthetic rollouts from a start state (Zhang et al., 2023; Treven et al., 2024). While this method significantly improves sample efficiency and data diversity, it suffers from compounding errors and fails to control the quality of generated trajectories. Other research leverages generative models to capture the distributions of collected datasets and randomly sample new transition data (Lu et al., 2024). Although these methods demonstrate some performance improvements, they remain inefficient when dealing with datasets containing limited optimal demonstrations. This inefficiency stems from their inability to effectively control the quality of generated data.
@@
We attempt to address this challenge by taking advantage of generative modelling to augment higher-quality data with directional goals. Unlike previous studies (Lu et al., 2024) that sample data unconditionally and randomly, we introduce GODA to incorporate representative goals, guiding the samples toward higher returns. Given the exceptional performance of diffusion models (Ho et al., 2020; Karras et al., 2022) in the field of generative artificial intelligence, GODA utilizes a diffusion model as its generative framework. GODA is trained to capture a comprehensive representation of the data distribution from the original dataset while sampling new data conditioned on selective high-return goals. This approach maximizes the utility of the limited well-performed trajectories in the original datasets.
@@
Inspired by Decision Transformer (Chen et al., 2021), we define the ‘goal’ as the return-to-go (RTG), which represents the cumulative rewards from the current step until the end, coupled with its specific timestep in trajectories. RTG explicitly indicates the expected future rewards for a given behavior at the current timestep for a specific trajectory. We assume that at the same timestep across different trajectories, a higher RTG signifies a higher goal. To generate samples that exceed the quality of the original dataset, we introduce three goal selection mechanisms and a scaling technique to control our expected goals during sampling.
@@
3) We show GODA’s competence through comprehensive experiments on the D4RL benchmark compared with state-of-the-art data augmentation methods across multiple offline RL algorithms. We further evaluate the effectiveness of GODA on a real-world application, i.e., traffic signal control, with small-size datasets obtained from widely used controllers in real-world deployments. These evaluations verify GODA’s effectiveness in addressing various challenges, significantly enhancing the applicability of RL-based methods for real-world scenarios.
@@
The remaining sections are organized as follows: Some preliminaries about offline RL and diffusion models are introduced in Section 2, followed by details about our methodology in Section 3. Section 4 describes the experimental settings about D4RL and TSC tasks, baselines and evaluation algorithms. Next, we show the data quality measurement for our augmented datasets in Section 5 and provide more detailed experimental results in Section 6. Finally, we conclude our research and discuss future directions in Section 7. The Related Work section is in Appendix A.
2 Preliminaries
2.1 Offline Reinforcement Learning
@@
In RL, the task environment is generally formulated as a Markov decision process (MDP) $\{\mathcal{S},\mathcal{A},\mathcal{R},\mathcal{P},\gamma\}$ (Sutton & Barto, 2018). $s\in\mathcal{S}$, $s^{\prime}\in\mathcal{S}$, $a\in\mathcal{A}$, $r=\mathcal{R}(s,a)$, $\mathcal{P}(s^{\prime}|s,a)$, and $\gamma\in[0,1)$ represent state, next state, action, reward function, state transition, and discount factor, respectively. RL aims to train an agent to interact with the environment and learn a policy $\pi$ from experience. The objective of RL is to maximize the expected discounted cumulative rewards over time:
@@
2.2 Diffusion Models
Diffusion models (Sohl-Dickstein et al., 2015; Ho et al., 2020; Karras et al., 2022), a class of well-known generative modeling methods, aim to learn a comprehensive representation of the data distribution $p_{\text{data}}(\mathbf{x}^{N})$ with a standard deviation $\sigma_{\text{data}}$ from a given dataset. Diffusion models generally have two primary processes, the forward process, also known as the diffusion process, and the reverse/sampling process.
@@ 2.2 Diffusion Models
For reverse process, Denoising Diffusion Probabilistic Model (DDPM) (Ho et al., 2020) models it as a Markov chain that involves denoising an initial noise $p\left(\mathbf{x}^{0}\right)=\mathcal{N}\left(\mathbf{x}^{0};\mathbf{0},\mathbf{I}\right)$ to the original data distribution with learned Gaussian transitions. Elucidated Diffusion Model (EDM) (Karras et al., 2022) formulates the forward and reverse processes as a probability-flow ordinary differential equation (ODE), where the noise level can be increased or decreased by moving the ODE forward or backward in time:
@@
where $\dot{\sigma}(t_{i})$ denotes derivative over denoise time and $\nabla_{\mathbf{x}}\log p(\mathbf{x};\sigma(t_{i}))$ is referred to as the score function (Song et al., 2020), which points towards regions of higher data density. It is worth noting that we use $t_{i}$ to denote the noise time to distinguish it from the trajectory timestep $t$. The ODE pushes the samples away from the data or closer to the data through infinitesimal forward or backward steps. The corresponding step sequence is $\left\{t_{0},t_{1},...t_{N}\right\}$, where $t_{N}=0$ and $N$ denotes the number of ODE solver iterations.
EDM proposes to estimate the score function using denoising score matching (Karras et al., 2022). Specifically, a denoiser neural network $D_{\theta}(\mathbf{x};\sigma)$ is trained to approximate data $\mathbf{x}^{N}$ sampled from $p_{\text{data}}$ by minimizing the $L_{2}$ denoising loss independently for each $\sigma$:
@@
Subsequently, the score function can be calculated as $\nabla_{\mathbf{x}}\log p(\mathbf{x};\sigma)=(D_{\theta}(\mathbf{x};\sigma)-\mathbf{x})/\sigma^{2}$. EDM employs Heun’s $2^{\text{nd}}$ order ODE solver (Ascher & Petzold, 1998) to solve the reverse-time ODE by iteratively refining $\mathbf{x}^{0}$ toward $t_{N}=0$, i.e., sampling data with the reverse process.
3 Methodology
Figure 2: Illustration of diffusion model training and data augmentation. The diffusion model is first trained on the original dataset to learn the underlying data distribution, and is then used to generate higher-quality samples conditioned on selected controllable goals.
In this section, we introduce GODA, a goal-conditioned data augmentation method utilizing generative modeling for augmenting higher-quality synthetic transition data. Our adopted diffusion model first learns comprehensive data distribution from the initial offline dataset, subsequently sampling new data towards higher return with controllable, selective goal conditions, as shown in Figure 2. In this part, we define a representative goal for GODA in Section 3.1 and introduce different selective mechanisms for goal conditions in Section 3.2. To proactively control the sampling direction, a controllable goal scaling factor is introduced in Section 3.3. For better integrating goal conditions as guidance within the diffusion model, we further propose a novel adaptive gated conditioning approach (Section 3.4) that introduces a condition-adaptive gate mechanism into the long skip connection and the residual connection. Section 3.5 presents the pseudocode (Algorithm 1) and outlines the implementation details of our method.
3.1 Return-oriented goal
Prior diffusion-based data augmentation work (Lu et al., 2024) lacks the ability to guide the sampling process toward a desired outcome. In our approach, we introduce a return-oriented goal as a conditioning signal for the diffusion model, analogous to how prompts are used to guide image generation in some diffusion models (Peebles & Xie, 2023; Li et al., 2025). As shown in Figure 3, the denoiser neural network not only takes as input the noised sample and noise level, as described in Equation 4, but also an extra goal condition. Inspired by Decision Transformer, we adopt return-to-go (RTG) (Chen et al., 2021), which represents the cumulative rewards from the current step till the end, as an explicit goal condition
@@ 3.1 Return-oriented goal
For each transition sample represented as a tuple $(s,a,s^{\prime},r)$ within a trajectory, the RTG serves as an unbiased measure of the future rewards corresponding to the current state-action pair. In the context of Decision Transformer, a higher RTG at a given timestep implies a more ambitious goal for the policy to achieve, thereby guiding the behavior of the agent toward more rewarding outcomes. Since the same behavior at different timesteps often yields varying RTGs across different trajectories, we combine the RTG with its corresponding timestep in the trajectory as the condition for each specific transition sample. The timestep signal acts as a timestamp for each goal.
3.2 Selective goal conditions
To get the training goal conditions from the original dataset, we first organize offline samples into trajectories, compute the RTG $\hat{g}_{t}$ for each, and append timestep $t$ to every sample. These goals along with the transition samples $\left\{(\hat{g}_{t},t,s_{t},a_{t},s_{t+1},r_{t})\right\}$ are then used to train the diffusion model to capture the underlying data distribution. For data augmentation (sampling procedure), however, we aim to generate new samples conditioned on a selectively chosen batch of goals. To fully leverage well-performing samples and augment samples with higher returns, we propose three distinct condition selection mechanisms: return-prior, RTG-prior, and random goal conditions.
(1) Return-prior goal condition. In this approach, we rank all trajectories based on their return values and select the top $n$ trajectories. During the sampling procedure of the diffusion model, the RTG and timestep pairs $(\hat{g}_{t},t)$ from these top $n$ trajectories are selected as the sampling goal conditions. This method filters high-return trajectories from the initial offline datasets and reuses them to sample more well-optimized transitions. The following example illustrates how we select goal conditions from the top two trajectories. $\hat{g}_{t}^{\tau}$ denotes the RTG value at timestep $t$ from trajectory $\tau$.
Original Trajectory $t=1$ $t=2$ $t=3$ $t=4$ $\cdots$ Return
1 (selected) $\hat{g}_{1}^{1}$ $\hat{g}_{2}^{1}$ $\hat{g}_{3}^{1}$ $\hat{g}_{4}^{1}$ $\cdots$ 100
2 (selected) $\hat{g}_{1}^{2}$ $\hat{g}_{2}^{2}$ $\hat{g}_{3}^{2}$ $\hat{g}_{4}^{2}$ $\cdots$ 95
3 (excluded) $\hat{g}_{1}^{3}$ $\hat{g}_{2}^{3}$ $\hat{g}_{3}^{3}$ $\hat{g}_{4}^{3}$ $\cdots$ 90
4 (excluded) $\hat{g}_{1}^{4}$ $\hat{g}_{2}^{4}$ $\hat{g}_{3}^{4}$ $\hat{g}_{4}^{4}$ $\cdots$ 70
(2) RTG-prior goal condition. We group RTGs by their associated timesteps and then sort them to select the top $n$ RTGs along with their corresponding timesteps as goal conditions. As illustrated in the example below, for each column, the RTG values are sorted in descending order from top to bottom. This approach selectively reuses high-RTG transitions for data augmentation, focusing on transitions that are most likely to yield higher returns.
Re-ordered Trajectory $t=1$ $t=2$ $t=3$ $t=4$ $\cdots$
1 (selected) $\hat{g}_{1}^{1}$ $\hat{g}_{2}^{1}$ $\hat{g}_{3}^{2}$ $\hat{g}_{4}^{4}$ $\cdots$
2 (selected) $\hat{g}_{1}^{2}$ $\hat{g}_{2}^{3}$ $\hat{g}_{3}^{1}$ $\hat{g}_{4}^{3}$ $\cdots$
3 (excluded) $\hat{g}_{1}^{3}$ $\hat{g}_{2}^{2}$ $\hat{g}_{3}^{3}$ $\hat{g}_{4}^{1}$ $\cdots$
4 (excluded) $\hat{g}_{1}^{4}$ $\hat{g}_{2}^{4}$ $\hat{g}_{3}^{4}$ $\hat{g}_{4}^{2}$ $\cdots$